Which is the eighth term in the sequence: a(base)n= - 1/25 (5^(n-1))

A. -3125
B. -125
C. -625
D. -25

Answer by stanbon(75887)

Which is the eighth term in the sequence: a(base)n= - 1/25 (5^(n-1))
a(8) = (1/25)(5^7) = 5^7/5^2 = 5^5 = 3125
============
Cheers,
Stan H.
----------
A. -3125
B. -125
C. -625
D. -25